OpenStudy (anonymous):

help please !!!! Find the next two terms in each sequence. 1. 12, 17, 22, 27, 32, . . .

OpenStudy (anonymous):

whats the change from one term to the next ?

OpenStudy (austinl):

\(12+\underline{~~~~}=17\\ 17+\underline{~~~~}=22\\ 22+\underline{~~~~}=27\)

OpenStudy (anonymous):

17-12= 22-17= 22-27=

OpenStudy (anonymous):

do you know the number yet its really easy i promise